given the formula, solve for the indicated variable
f=9/5c+32 for c

Answer by jim_thompson5910(35256)

f=9/5c+32


f-32=9/5c


(5/9)(f-32)=c


c = (5/9)(f-32)


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim